OpenStudy (anonymous):

I have a question on rational exponents. It's 24/4(3/2).

OpenStudy (whpalmer4):

You should use the Equation button to enter things like this, so we don't have to guess what the formula really is... I'll assume that it is \[\frac{24}{4^{(3/2)}}\] Remember that \[x^{(a/b)} = \sqrt[b]{x^a}\] so your denominator could be written as \[\sqrt{4*4*4}\] which should help you along...